OpenStudy (anonymous):

Graph the equation -5-x= 0.

OpenStudy (skullpatrol):

Use the addition property of equality and add x to both sides please.

OpenStudy (skullpatrol):

$$-5-x= 0$$ $$-5-x+x= 0+x$$

OpenStudy (skullpatrol):

Simplify please.

OpenStudy (anonymous):

ok...i will try..

OpenStudy (anonymous):

i don't know.. :(

OpenStudy (skullpatrol):

What does $$-x+x$$equal?